# Gradient Interference-Aware Graph Coloring for Multitask Learning

### v1

@@ F.1 Setting, definitions, and population objects
$\mu_{i}\;\in\;\mathbb{R}^{d}\qquad(i=1,\dots,K)$ (F.1)
@@
$C^{\star}_{ij}\;=\;\frac{\langle\mu_{i},\mu_{j}\rangle}{\|\mu_{i}\|\,\|\mu_{j}\|},\qquad i\neq j,\quad C^{\star}_{ii}=1.$ (F.2)
Definition F.2.
@@
F.2 Assumptions
We adopt the following assumptions, which are standard in analyses of stochastic-gradient methods and verifiable in practice (see, e.g., [79, 80, 81, 82, 83], for concentration of geometrically weighted and mixing sequences, see [84, 85]).
@@ F.2 Assumptions
$\tilde{g}_{i}\;=\;\sum_{t=1}^{R}w_{t}\,g_{i,t},\qquad w_{t}\;=\;\frac{(1-\beta)\,\beta^{R-t}}{1-\beta^{R}},\qquad\beta\in[0,1).$ (F.5)
@@
$n_{\rm eff}^{-1}\;:=\;\sum_{t=1}^{R}w_{t}^{2}\quad\text{so that}\quad c_{1}\,\frac{1-\beta^{2}}{1-\beta^{2R}}\;\leq\;\frac{1}{n_{\rm eff}}\;\leq\;c_{2}\,\frac{1-\beta^{2}}{1-\beta^{2R}}$ (F.6)
for absolute constants $0<c_{1}\leq c_{2}<\infty$ (in particular $n_{\rm eff}=\Theta\!\big(\tfrac{1-\beta^{2R}}{1-\beta^{2}}\big)$).
@@
F.3 Deterministic group recovery from the conflict graph
We begin with basic graph-thoeretic facts that we will use once we have established that the empirical conflict graph coincides with its population counterpart.
@@ F.3 Deterministic group recovery from the conflict graph
F.4 Uniform control of empirical cosines from EMA gradients
@@
$\Pr\!\Big(\,\big|\langle\tilde{g}_{i}-\mu_{i},\ u\rangle\big|>\varepsilon\,\Big)\ \leq\ 2\exp\!\Big(-c\,n_{\rm eff}\,\varepsilon^{2}/\sigma^{2}\Big).$ (F.7)
@@
$\Pr\!\Big(\,\max_{1\leq j\leq M}\big|\langle\tilde{g}_{i}-\mu_{i},\ u_{j}\rangle\big|>\varepsilon\,\Big)\ \leq\ 2\,M\,\exp\!\Big(-c\,n_{\rm eff}\,\varepsilon^{2}/\sigma^{2}\Big).$ (F.8)
@@
$\big|\langle\tilde{g}_{i}-\mu_{i},\ \tfrac{\mu_{j}}{\|\mu_{j}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{j}-\mu_{j},\ \tfrac{\mu_{i}}{\|\mu_{i}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{i}-\mu_{i},\ \tfrac{\mu_{i}}{\|\mu_{i}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{j}-\mu_{j},\ \tfrac{\mu_{j}}{\|\mu_{j}\|}\rangle\big|\leq\epsilon,$ (F.9)
@@
$\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \frac{6\,\epsilon}{m_{0}}\,+\,\frac{4\,\epsilon^{2}}{m_{0}^{2}}.$ (F.10)
@@
$\langle\tilde{g}_{i},\tilde{g}_{j}\rangle-\langle\mu_{i},\mu_{j}\rangle=\langle\delta_{i},\mu_{j}\rangle+\langle\mu_{i},\delta_{j}\rangle+\langle\delta_{i},\delta_{j}\rangle,$ (F.11)
@@
$\|\tilde{g}_{i}\|=\|\mu_{i}\|\sqrt{1+2\langle\delta_{i},\mu_{i}\rangle/\|\mu_{i}\|^{2}+\|\delta_{i}\|^{2}/\|\mu_{i}\|^{2}}$ (F.12)
@@
$|\langle\delta_{i},\mu_{j}/\|\mu_{j}\|\rangle|\leq\epsilon$ (F.13)
@@
$|\langle\delta_{i},\mu_{i}/\|\mu_{i}\|\rangle|\leq\epsilon$ (F.14)
@@
$|\langle\delta_{i},\mu_{j}\rangle|\leq\epsilon\|\mu_{j}\|$ (F.15)
@@
$|\langle\delta_{i},\mu_{i}\rangle|\leq\epsilon\|\mu_{i}\|$ (F.16)
@@
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\epsilon^{2}}\,\log\!\Big(\frac{K}{\delta}\Big)$ (F.17)
@@
$\max_{i<j}\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \epsilon$ (F.18)
@@
F.5 Exact edge recovery and group recovery
@@
$\max_{i<j}|\widehat{C}_{ij}-C^{\star}_{ij}|\leq\epsilon$ (F.19)
with $\epsilon<\gamma$
$\max_{i<j}|\widehat{C}_{ij}-C^{\star}_{ij}|\leq\epsilon\qquad\text{with}\qquad\epsilon<\gamma$ (F.20)
then
$\widehat{C}_{ij}\geq-(\tau-\gamma)-\epsilon>-\tau$ (F.21)
@@
Under Assumptions 15, if
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Big(\frac{K}{\delta}\Big)$ (F.22)
@@
We now translate exact edge recovery into group recovery.
@@
F.6 Quantitative probe-budget requirement
@@
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Big(\frac{K}{\delta}\Big)\qquad\Big(\text{equivalently,}\quad\sum_{t=1}^{R}w_{t}^{2}\ \leq\ c\,\frac{m_{0}^{2}\,\gamma^{2}}{\sigma^{2}}\,\frac{1}{\log(K/\delta)}\Big)$ (F.23)
then $\Pr(\widehat{G}=G^{\star})\geq 1-\delta$, and consequently Theorems 67 apply. In particular, at fixed $\beta$ and for large $R$, $n_{\rm eff}=\Theta((1-\beta^{2})^{-1})$, recovering the usual $O(\log K)$ dependence on the number of tasks.
F.7 Summary of the recovery argument
@@
    $\max_{i<j}\bigl|\widehat{C}_{ij}-C^{\star}_{ij}\bigr|\ \leq\ \epsilon,$ (F.24)
@@
$n_{\mathrm{eff}}\;\geq\;C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Bigl(\frac{K^{2}}{\delta}\Bigr)$ (F.25)
@@
Appendix G Computational Complexity of One Refresh (and Amortized Over Training)

### v2

@@ B.1 Setting, definitions, and population objects
$\mu_{i}\;\in\;\mathbb{R}^{d}\qquad(i=1,\dots,K)$ (14)
@@
$C^{\star}_{ij}\;=\;\frac{\langle\mu_{i},\mu_{j}\rangle}{\|\mu_{i}\|\,\|\mu_{j}\|},\qquad i\neq j,\quad C^{\star}_{ii}=1.$ (15)
Definition B.2.
@@
B.2 Assumptions
We adopt the following assumptions, which are standard in analyses of stochastic-gradient methods and verifiable in practice (see, e.g., Robbins & Monro 1951; Kushner & Yin 2003; Nemirovski et al. 2009; Bottou et al. 2018; Wainwright 2019; for concentration of geometrically weighted and mixing sequences, see Merlevède et al. 2011; De la Pena et al. 2009).
@@ B.2 Assumptions
$\tilde{g}_{i}\;=\;\sum_{t=1}^{R}w_{t}\,g_{i,t},\qquad w_{t}\;=\;\frac{(1-\beta)\,\beta^{R-t}}{1-\beta^{R}},\qquad\beta\in[0,1).$ (18)
@@
$n_{\rm eff}^{-1}\;:=\;\sum_{t=1}^{R}w_{t}^{2}\;=\;\frac{(1-\beta)^{2}(1-\beta^{2R})}{(1-\beta^{R})^{2}(1-\beta^{2})}.$ (19)
In particular, as $R\to\infty$ (with fixed $\beta\in[0,1)$), we have $n_{\rm eff}\to\frac{1+\beta}{1-\beta}$.
@@
B.3 Deterministic group recovery from the conflict graph
We begin with basic graph-theoretic facts that we will use once we have established that the empirical conflict graph coincides with its population counterpart.
@@ B.3 Deterministic group recovery from the conflict graph
B.4 Uniform control of empirical cosines from EMA gradients
@@
$\Pr\!\Big(\,\big|\langle\tilde{g}_{i}-\mu_{i},\ u\rangle\big|>\varepsilon\,\Big)\ \leq\ 2\exp\!\Big(-c\,n_{\rm eff}\,\varepsilon^{2}/\sigma^{2}\Big).$ (20)
@@
$\Pr\!\Big(\,\max_{1\leq j\leq M}\big|\langle\tilde{g}_{i}-\mu_{i},\ u_{j}\rangle\big|>\varepsilon\,\Big)\ \leq\ 2\,M\,\exp\!\Big(-c\,n_{\rm eff}\,\varepsilon^{2}/\sigma^{2}\Big).$ (21)
@@
$\big|\langle\tilde{g}_{i}-\mu_{i},\ \tfrac{\mu_{j}}{\|\mu_{j}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{j}-\mu_{j},\ \tfrac{\mu_{i}}{\|\mu_{i}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{i}-\mu_{i},\ \tfrac{\mu_{i}}{\|\mu_{i}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{j}-\mu_{j},\ \tfrac{\mu_{j}}{\|\mu_{j}\|}\rangle\big|\leq\epsilon,$ (22)
@@
$\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \frac{6\,\epsilon}{m_{0}}\,+\,\frac{4\,\epsilon^{2}}{m_{0}^{2}}.$ (23)
@@
$\langle\tilde{g}_{i},\tilde{g}_{j}\rangle-\langle\mu_{i},\mu_{j}\rangle=\langle\delta_{i},\mu_{j}\rangle+\langle\mu_{i},\delta_{j}\rangle+\langle\delta_{i},\delta_{j}\rangle,$ (24)
@@
$\|\tilde{g}_{i}\|=\|\mu_{i}\|\sqrt{1+2\langle\delta_{i},\mu_{i}\rangle/\|\mu_{i}\|^{2}+\|\delta_{i}\|^{2}/\|\mu_{i}\|^{2}}$ (25)
@@
$|\langle\delta_{i},\mu_{j}/\|\mu_{j}\|\rangle|\leq\epsilon$ (26)
@@
$|\langle\delta_{i},\mu_{i}/\|\mu_{i}\|\rangle|\leq\epsilon$ (27)
@@
$|\langle\delta_{i},\mu_{j}\rangle|\leq\epsilon\|\mu_{j}\|$ (28)
@@
$|\langle\delta_{i},\mu_{i}\rangle|\leq\epsilon\|\mu_{i}\|$ (29)
@@
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\epsilon^{2}}\,$ (30)
@@
$\max_{i<j}\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \epsilon$ (31)
@@
B.5 Exact edge recovery and group recovery
@@
$\max_{i<j}\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \epsilon\qquad\text{with}\qquad\epsilon<\gamma,$ (32)
then the empirical conflict graph equals the population graph:
$\widehat{G}\;=\;G^{\star}.$ (33)
Equivalently, for every $i\neq j$,
$C^{\star}_{ij}\leq-(\tau+\gamma)\ \Rightarrow\ \widehat{C}_{ij}<-\tau\quad\text{and}\quad C^{\star}_{ij}\geq-(\tau-\gamma)\ \Rightarrow\ \widehat{C}_{ij}>-\tau.$ (34)
@@
Under Assumptions 15, there exists a universal constant $C>0$ such that if
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Big(\frac{K^{2}}{\delta}\Big),$ (35)
@@
B.6 Quantitative probe-budget requirement
@@
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\qquad\Big(\text{equivalently,}\quad\sum_{t=1}^{R}w_{t}^{2}\ \leq\ c\,\frac{m_{0}^{2}\,\gamma^{2}}{\sigma^{2}}\,\frac{1}{\log(K/\delta)}\Big)$ (36)
then $\Pr(\widehat{G}=G^{\star})\geq 1-\delta$, and consequently Theorems 34 apply. In particular, for fixed $\beta$ and large $R$, $n_{\rm eff}\to\frac{1+\beta}{1-\beta}$ (i.e., it saturates). Thus, to meet the required budget as $K$ grows, one increases $n_{\rm eff}$ by choosing $\beta$ closer to 11 (e.g., $1-\beta\asymp 1/\log(K^{2}/\delta)$), or by switching to a unnormalized averaging approach.
B.7 Summary of the recovery argument
@@
    $\max_{i<j}\bigl|\widehat{C}_{ij}-C^{\star}_{ij}\bigr|\ \leq\ \epsilon,$ (37)
@@
$n_{\mathrm{eff}}\;\geq\;C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Bigl(\frac{K^{2}}{\delta}\Bigr)$ (39)
@@
Appendix C Descent Bounds for Scheduled versus Aggregated Updates
$F(y)\ \leq\ F(x)+\langle\nabla F(x),y-x\rangle+\tfrac{L}{2}|y-x|^{2},\quad\forall x,y.$ (40)
C.1 Single refresh baselines and notation
C.1.1 Single aggregated step
Definition C.1 (Aggregated step).
$x^{\mathrm{agg}}\ :=\ x-\eta\,\sum_{r=1}^{m}G_{r}^{0}.$ (41)
One-shot $L$-smoothness bound.
Applying $L$-smoothness with $y=x^{\mathrm{agg}}$ yields
$F(x^{\mathrm{agg}})\ \leq\ F(x)\;-\;\eta\Big\langle\nabla F(x),\sum_{r=1}^{m}G_{r}^{0}\Big\rangle\;+\;\tfrac{L\eta^{2}}{2}\Big\|\sum_{r=1}^{m}G_{r}^{0}\Big\|^{2}.$ (42)
C.1.2 Scheduled group sequence over one refresh
Definition C.2 (Scheduled refresh).
Starting from the same point $x$, define
$x_{0}:=x,\qquad x_{r}:=x_{r-1}-\eta\,G_{r}(x_{r-1})\quad(r=1,\ldots,m),\qquad x^{\mathrm{sch}}:=x_{m}.$ (43)
Order and notation.
C.2 Telescoping bound for scheduled updates
Lemma 3 (Smoothness Expansion for Two Scheduled Groups).
Let $m=2$ and $G_{r}^{0}:=G_{r}(x)$. For any $\eta\in(0,1/L]$,
$F(x^{\text{sch}})$ $\leq F(x)\;-\;\eta\,\langle\nabla F(x),\,G_{1}^{0}\rangle\;+\;\tfrac{L\eta^{2}}{2}\,\|G_{1}^{0}\|^{2}$ (44)
Proof sketch.
$\nabla F(x_{1})=\nabla F(x)-\int_{0}^{1}H(x{-}t\eta G_{1}^{0})\,\eta G_{1}^{0}\,dt$ (45)
by the fundamental theorem of calculus along the segment $x\to x_{1}$. ∎
C.2.1 Start-of-refresh reduction under per-group lipschitzness
We adopt the following assumption whenever we compare bounds solely in terms of start-of-refresh measurements. It will be used throughout Sections C.3C.6
Assumption 6 (Per-group lipschitzness).
Each group map $G_{r}(\cdot)$ is $L_{r}$-lipschitz:
$\|G_{r}(u)-G_{r}(v)\|\ \leq\ L_{r}\,\|u-v\|\quad\text{for all }u,v.$ (46)
$\|G_{2}(x_{1})\|\leq\|G_{2}^{0}\|+L_{2}\eta\|G_{1}^{0}\|$ (47)
For general $m$
$\|G_{r}(x_{r-1})\|\ \leq\ \|G_{r}^{0}\|+L_{r}\,\eta\sum_{p<r}\|G_{p}^{0}\|\qquad(r=2,\dots,m)$ (48)
C.3 Upper bounds for scheduled and aggregated updates (general $m$)
Applying $L$-smoothness $m$ times yields the scheduled upper bound
$\mathrm{UB}_{\text{sch}}(x;\eta)$ $=F(x)\;-\;\eta\,\sum_{r=1}^{m}\big\langle\nabla F(x),\,G_{r}(x_{r-1})\big\rangle\;+\;\frac{L\eta^{2}}{2}\,\sum_{r=1}^{m}\|G_{r}(x_{r-1})\|^{2}$ (49)
The aggregated upper bound is the one-shot bound from Equation 42, restated as
$\mathrm{UB}_{\text{agg}}(x;\eta):=F(x)\;-\;\eta\,\Big\langle\nabla F(x),\sum_{r=1}^{m}G_{r}^{0}\Big\rangle\;+\;\frac{L\eta^{2}}{2}\,\Big\|\sum_{r=1}^{m}G_{r}^{0}\Big\|^{2}$ (50)
The integrals in Equation 49 are over ordered pairs $p<q$ along the specific sequence $x_{0}\to x_{1}\to\cdots\to x_{m}$; the bound therefore depends on the within-refresh order. Randomizing the order yields an expected version.
In Sections C.4C.6 we express the scheduled bound in terms of $\{G_{r}^{0}\}$ under the per-group lipschitzness assumption. The associated drift terms are aggregated into $R_{m}(x;\eta)$.
C.4 Scheduled and aggregated gap at a common linearization
Define the shorthand
$I_{pq}(x;\eta)\ :=\ \int_{0}^{1}\!\big\langle H(x{-}t\eta G_{p}^{0})\,G_{p}^{0},\ G_{q}^{0}\big\rangle\,dt$ (51)
Theorem 5 (Upper-bound gap under per-group lipschitzness).
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\;\leq\;\eta^{2}\!\sum_{1\leq p<q\leq m}\!\Big(\!-L\,\langle G_{p}^{0},G_{q}^{0}\rangle\;+\;I_{pq}(x;\eta)\Big)\;+\;R_{m}(x;\eta).$ (52)
Using $\|H(\cdot)\|_{\mathrm{op}}\leq L$ and Cauchy-Schwarz (Steele, 2004)
$I_{pq}(x;\eta)\ \leq\ L\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|$ (53)
which gives the envelope
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\ \leq\ L\eta^{2}\!\sum_{p<q}\!\big(\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|-\langle G_{p}^{0},G_{q}^{0}\rangle\big)\;+\;R_{m}(x;\eta)\ \ \geq\ 0$ (54)
Interpretation
Proposition 4 (Drift penalty bound under per-group lipschitzness).
Assume each group map $G_{r}$ is $L_{r}$-lipschitz. Then for $r\geq 2$,
$\|G_{r}(x_{r-1})\|\ \leq\ \|G_{r}^{0}\|+L_{r}\eta\sum_{p<r}\|G_{p}^{0}\|\ :=\ \|G_{r}^{0}\|+L_{r}\eta\,S_{r-1},$ (55)
and the scheduled start substitution error satisfies
$R_{m}(x;\eta)\ \leq$ $\ \eta^{2}\Big(\sum_{p=1}^{m}\|G_{p}^{0}\|\Big)\,\sum_{r=2}^{m}L_{r}\,S_{r-1}$ (56)
C.5 Sufficient conditions for a tighter scheduled bound
Assumption 7 (Hessian-weighted negative cross-terms).
There exist nonnegative margins $\{\Gamma_{pq}\}_{p<q}$ such that
$I_{pq}(x;\eta)\ =\ \int_{0}^{1}\!\big\langle H(x{-}t\eta G_{p}^{0})\,G_{p}^{0},\ G_{q}^{0}\big\rangle dt\ \leq\ -\,\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|\quad\text{for all }p<q$ (57)
Assuming per-group lipschitzness and 57, for any $\eta\in(0,1/L]$,
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\ \leq\ \eta^{2}\!\sum_{p<q}\!\Big(\!-L\,\langle G_{p}^{0},G_{q}^{0}\rangle\;-\;\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|\Big)\ +\ R_{m}(x;\eta)$ (58)
In particular, if
$\sum_{p<q}\!\Big(\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|+L\,\langle G_{p}^{0},G_{q}^{0}\rangle\Big)\ >\ \frac{R_{m}(x;\eta)}{\eta^{2}}$ (59)
then $\mathrm{UB}_{\text{sch}}(x;\eta)\ <\ \mathrm{UB}_{\text{agg}}(x;\eta)$
C.6 PL or strong convexity: standard rate and upper-bound gains for scheduling
$\tfrac{1}{2}|\nabla F(x)|^{2}\ \geq\ \mu\,(F(x)-F^{\star}),\qquad\forall x$ (60)
$F(x^{\text{agg}})\ \leq\ F(x)\;-\;\eta\Big(1-\tfrac{L\eta}{2}\Big)\,|\nabla F(x)|^{2}\ \leq\ \Big(1-2\mu\eta\big(1-\tfrac{L\eta}{2}\big)\Big)\,\big(F(x)-F^{\star}\big)$ (61)
$\Delta{\text{UB}}(x;\eta)\ :=\ \mathrm{UB}_{\text{agg}}(x;\eta)-\mathrm{UB}_{\text{sch}}(x;\eta)\ \geq\ 0$ (62)
whenever 59 holds. Since $F(x^{\text{sch}})\leq\mathrm{UB}_{\text{sch}}(x;\eta)$ and $\mathrm{UB}_{\text{agg}}(x;\eta)$ upper-bounds the one-shot decrease term in 61, we obtain the bound-level contraction
$F(x^{\text{sch}})-F^{\star}\ \leq\ \Big(1-2\mu\eta\big(1-\tfrac{L\eta}{2}\big)\Big)\,\big(F(x)-F^{\star}\big)\;-\;\Delta_{\text{UB}}(x;\eta).$ (63)
Consequently, under per-group lipschitzness and 59, the scheduled refresh satisfies the standard gradient-descent contraction and, in addition, achieves an extra nonnegative decrement $\Delta_{\mathrm{UB}}(x;\eta)$ in the upper bound.
C.7 Why the assumptions are mild
C.7.1 $L$-smoothness
This is the same regularity used throughout the main paper and in our baselines. Each task loss we optimize is $L_{i}$-smooth, so the overall objective is $L$-smooth. We only use this to apply the standard smoothness (descent) inequality (Nesterov, 2004; Beck, 2017).
C.7.2 Per-group lipschitzness of $G_{r}$
C.7.3 Negative Hessian-weighted cross-terms
C.7.4 PL and strong convexity
We invoke PL only to convert a per-refresh decrease into a standard contraction factor. We do not require global strong convexity. A local PL inequality around the iterates is enough, which is commonly observed after warm-up and annealing we already use (Karimi et al., 2016; Zhou et al., 2021; Liu, 2025).
C.8 Concluding remarks
This appendix formalizes a bound-level comparison between scheduled and aggregated updates. Without additional structure the scheduled bound need not be tighter, but under per-group lipschitzness and negative Hessian-weighted cross-terms it becomes strictly tighter, and under PL the scheduled refresh inherits the standard GD contraction with an additional nonnegative decrement. In practice, these conditions arise naturally once the task-group assignments stabilize, so the scheduler will typically achieve tighter descent bounds without changing step sizes or gradient information.
Appendix D Computational Complexity of One Refresh (and Amortized Over Training)